On January 1, We Grow Awesome Grass issued and sold a $350,000, 6%, 10-year bond payable, and received proceeds of $375,000. Interest is payable each June 30 and December 31. The company uses the straight-line method to amortize the premium. The carrying value of the bonds immediately after the first interest payment is:

$376,250.

$351,250.

$348,750.

$373,750.

$350,000.
Answer #1

Premium on bonds payable = 375000-350000 = 25000

Semiannual amortization = 25000/20 = 1250

Carrying value = 375000-1250 = 373750

So answer is d) $373750